# On the Sequence Evaluation based on Stochastic Processes

### v3

@@ A.1 Proof of Proposition 1
Given that $s_{t}-\mu_{t}={\mathbf{W}}(B_{1}(t),\ldots,B_{d}(t))^{\top}$, and considering the independence of $B_{1}(t),\ldots,B_{d}(t)$ along with the properties of the standard BB, we have for any $t,t^{\prime}\in\{1,2,\ldots,T-1\}$: $\mathrm{E}[s_{t}-\mu_{t}]=0$, $\mathrm{Var}[s_{t}]=[\Sigma_{T}]_{t,t}\Sigma$ and ${\rm Cov}[s_{t},s_{t^{\prime}}]=[\Sigma_{T}]_{t,t^{\prime}}\Sigma$. Therefore, the vectorized form of ${\mathbf{s}}-\bm{\mu}$ follows a multivariate normal distribution:
$\mathrm{vec}({\mathbf{s}}-\bm{\mu})\sim N(0,\Sigma_{T}\otimes\Sigma),$
@@
$L(\Sigma|\bar{{\mathbf{s}}})=(2\pi)^{-d(T-1)/2}\lvert\Sigma_{T}\otimes\Sigma\rvert^{-1/2}\exp[-\mathrm{vec}({\mathbf{s}}-\bm{\mu})^{\top}[\Sigma_{T}\otimes\Sigma]^{-1}\mathrm{vec}({\mathbf{s}}-\bm{\mu})/2].$
@@
$\mathrm{vec}({\mathbf{s}}-\bm{\mu})^{\top}[\Sigma_{T}\otimes\Sigma]^{-1}\mathrm{vec}({\mathbf{s}}-\bm{\mu})=\mathrm{vec}({\mathbf{s}}-\bm{\mu})^{\top}[\Sigma_{T}^{-1}\otimes\Sigma^{-1}]\mathrm{vec}({\mathbf{s}}-\bm{\mu})$
$=$ $\mathrm{vec}({\mathbf{s}}-\bm{\mu})^{\top}\mathrm{vec}(\Sigma^{-1}({\mathbf{s}}-\bm{\mu})\Sigma_{T}^{-1})=\mathrm{tr}(({\mathbf{s}}-\bm{\mu})^{\top}\Sigma^{-1}({\mathbf{s}}-\bm{\mu})\Sigma_{T}^{-1})$
$=$ $\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}-\bm{\mu})\Sigma_{T}^{-1}({\mathbf{s}}-\bm{\mu})^{\top}).$
@@
$L(\Sigma|\bar{{\mathbf{s}}})=(2\pi)^{-d(T-1)/2}\lvert\Sigma_{T}\rvert^{-d/2}\lvert\Sigma\rvert^{-(T-1)/2}\exp[-\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}-\bm{\mu})\Sigma_{T}^{-1}({\mathbf{s}}-\bm{\mu})^{\top})/2].$
@@
$\ell(\Sigma|\bar{{\mathbf{s}}})=-\frac{d(T-1)}{2}\log(2\pi)-\frac{d}{2}\log(\lvert\Sigma_{T}\rvert)-\frac{(T-1)}{2}\log(\lvert\Sigma\rvert)-\frac{1}{2}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}-\bm{\mu})\Sigma_{T}^{-1}({\mathbf{s}}-\bm{\mu})^{\top}).$
@@
$=$ $-\frac{d\sum_{i=1}^{n}(T_{i}-1)}{2}\log(2\pi)-\frac{d}{2}\sum_{i=1}^{n}\log(\lvert\Sigma_{T_{i}}\rvert)-\frac{\sum_{i=1}^{n}(T_{i}-1)}{2}\log(\lvert\Sigma\rvert)$
$-\frac{1}{2}\sum_{i=1}^{n}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}).$
A.2 Proof of Proposition 2
@@
To find the MLE of $\Sigma$, we need to minimize the negative log-likelihood function, which is equivalent to minimize
$g(\Sigma)=\sum_{i=1}^{n}(T_{i}-1)\log(\lvert\Sigma\rvert)+\sum_{i=1}^{n}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}).$
Since $\Sigma={\mathbf{W}}{\mathbf{W}}^{\top}$ is positive definite, we can compute the gradient of $g(\Sigma)$ with respect to $\Sigma$. Note that $(\mathrm{d}/\mathrm{d}\Sigma)\log(\lvert\Sigma\rvert)=(\Sigma^{\top})^{-1}=\Sigma^{-1}$ and
$\frac{\mathrm{d}}{\mathrm{d}\Sigma}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top})$
$=$ $-(\Sigma^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}\Sigma^{-1})^{\top}=-\Sigma^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}\Sigma^{-1}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}\Sigma}g(\Sigma)=\Big{(}\sum_{i=1}^{n}(T_{i}-1)\Big{)}\Sigma^{-1}-\Sigma^{-1}\Big{(}\sum_{i=1}^{n}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}\Big{)}\Sigma^{-1}.$
@@
$\widehat{\Sigma}=\Big{(}\sum_{i=1}^{n}(T_{i}-1)\Big{)}^{-1}\Big{(}\sum_{i=1}^{n}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}\Big{)}.$
Appendix B Training Details
B.1 SP Encoder
B.2 Hyper-parameter Tuning
While training the WikiSection SP Encoder, we experimented with different $\epsilon$ in $L_{\rm NLL}$ to see its impact on the performance of the trained encoder. Note that $\epsilon$ determines the perturbation added to the matrix $\widehat{\Sigma}$. The eigenvalues of the initial $\widehat{\Sigma}$ range from $10^{-6}$ to $10^{-1}$, with the majority of which lying in $[10^{-3},10^{-5}]$. Thus we tested the following three different $\epsilon$:
We choose the small $\epsilon$ based on the performance.
Appendix C Other scores used in this paper
@@
Barzilay & Lapata (2005) is the most recognized entity-based approach. It creates a two-way contingency table for each input document to track the appearance of entities in each sentence. We use Stanford’s CoreNLP to annotate the documents and the implementation provided in the Coheoka library [^2] https://github.com/kigawas/coheoka to obtain the Entity Grid score.
@@
Sheng et al. (2024) introduces BBScore, and also check the main text for a comprehensive comparison between BBScore and SPM.
SPM + CLS
We use SPM and local moving window SPM with window sizes being 1, 2, 3, 5, and 10 as features showing how coherent the text is and an conceptual figure is illustrated in Figure 3 . A three-layer perceptron ($1\times 2048\times 512\times 1$ for global task and $1\times 512\times 1024\times 1$ for local task) then takes those features as input and outputs predictions for designed tasks.
Figure 3: Illustration of SPM over the entire trajectory and SPM with a local moving window (size=2).
Appendix D BBScore and SPM under different article length
As illustrated in Fig 4, we present a visualization of both the BBScore and SPM for the original and block-shuffled scores. From the plot, it is evident that the SPM remains more stable, regardless of the article length. Notably, using SPM, the shuffled scores (in red) exhibit a similar pattern to the unshuffled scores under different article length.
Figure 4: Visualization of BBscore and SPM.
Appendix E Comparison of Latents encoded by CL and SP Encoder
First, we conduct a comparison between the BB encoder and the CL encoder. As shown in Fig 5 (A.1) and (B.1), we plot the rough shape of the latent trajectory (blue and red regions) and the target BB (green region), where we observe that the BB encoder appears to learn the target BB more accurately.
To further highlight the presence of ineffective dimensions in the CL Encoder, Figures 5 (A.2) and (B.2) display the latent trajectory of a sample article encoded by the CL Encoder, which reveals only two highly symmetric effective dimensions. In contrast, the SP Encoder captures a more diverse and nuanced trajectory, reflecting superior dimensional effectiveness. This observation may also elucidate why Wang et al. (2022) could not establish a clear relationship between encoder dimensions and the performance of the results.
Figure 5: CL Encoder vs SP Encoder in the latent representation: Latent analysis of the CL Encoder (A.1 and A.2) and SP Encoder (B.1 and B.2). (A.2) and (B.2) show an example of the latent representation of an article, with different colors representing different dimensions (total dimension=32).
Appendix F Difference between WikiSection and WikiText
[…]
Appendix G Human-AI comparison test
Table 6 presents the performance of SPM computed with different $\widehat{\Sigma}\in\mathbb{R}^{d}$, while Table 7 shows the performance of the BBScore with various $\widehat{\sigma}\in\mathbb{R}$, where the subscript indicates the dataset used for approximation.
@@ Appendix G Human-AI comparison test
Table 8: Human-AI Task Results with DetectGPT (Mitchell et al., 2023). As a comparison, SPM only requires one LLM inference.

### v4

@@ C.1 Proof of Proposition 1
Given that $s_{t}-\mu_{t}={\mathbf{W}}(B_{1}(t),\ldots,B_{d}(t))^{\top}$, and considering the independence of $B_{1}(t),\ldots,B_{d}(t)$ along with the properties of the standard BB, we have for any $t,t^{\prime}\in\{1,2,\ldots,T-1\}$: $\mathrm{E}[s_{t}-\mu_{t}]=0$, $\mathrm{Var}[s_{t}]=[\Sigma_{T}]_{t,t}\Sigma$ and ${\rm Cov}[s_{t},s_{t^{\prime}}]=[\Sigma_{T}]_{t,t^{\prime}}\Sigma$. Therefore, the vectorized form of ${\mathbf{s}}-\boldsymbol{\mu}$ follows a multivariate normal distribution:
$\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})\sim N(0,\Sigma_{T}\otimes\Sigma),$
@@
$L(\Sigma|\bar{{\mathbf{s}}})=$ $(2\pi)^{-d(T-1)/2}\lvert\Sigma_{T}\otimes\Sigma\rvert^{-1/2}$
$\cdot\exp[-\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})^{\top}[\Sigma_{T}\otimes\Sigma]^{-1}\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})/2].$
@@
$\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})^{\top}[\Sigma_{T}\otimes\Sigma]^{-1}\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})$
$=\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})^{\top}[\Sigma_{T}^{-1}\otimes\Sigma^{-1}]\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})$
$=\mathrm{vec}({\mathbf{s}}-\boldsymbol{\mu})^{\top}\mathrm{vec}(\Sigma^{-1}({\mathbf{s}}-\boldsymbol{\mu})\Sigma_{T}^{-1})$
$=\mathrm{tr}(({\mathbf{s}}-\boldsymbol{\mu})^{\top}\Sigma^{-1}({\mathbf{s}}-\boldsymbol{\mu})\Sigma_{T}^{-1})$
$=\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}-\boldsymbol{\mu})\Sigma_{T}^{-1}({\mathbf{s}}-\boldsymbol{\mu})^{\top}).$
@@
$L(\Sigma|\bar{{\mathbf{s}}})=$ $(2\pi)^{-d(T-1)/2}\lvert\Sigma_{T}\rvert^{-d/2}\lvert\Sigma\rvert^{-(T-1)/2}$
$\cdot\exp[-\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}-\boldsymbol{\mu})\Sigma_{T}^{-1}({\mathbf{s}}-\boldsymbol{\mu})^{\top})/2].$
@@
[…]
$-\frac{1}{2}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}-\boldsymbol{\mu})\Sigma_{T}^{-1}({\mathbf{s}}-\boldsymbol{\mu})^{\top}).$
@@
[…]
$-\frac{1}{2}\sum_{i=1}^{n}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})^{\top}).$
C.2 Proof of Proposition 2
@@
To find the MLE of $\Sigma$, we need to minimize the negative log-likelihood function, which is equivalent to minimizing:
$g(\Sigma)=\sum_{i=1}^{n}(T_{i}-1)\log(\lvert\Sigma\rvert)+\sum_{i=1}^{n}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})^{\top}).$ (8)
Since $\Sigma={\mathbf{W}}{\mathbf{W}}^{\top}$ is positive definite, we can compute the gradient of $g(\Sigma)$ with respect to $\Sigma$. Note that:
$\frac{\mathrm{d}}{\mathrm{d}\Sigma}\log(\lvert\Sigma\rvert)=\Sigma^{-1},$
$\frac{\mathrm{d}}{\mathrm{d}\Sigma}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})^{\top})$
$=$ $-\Sigma^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})^{\top}\Sigma^{-1}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}\Sigma}g(\Sigma)=\Big(\sum_{i=1}^{n}(T_{i}-1)\Big)\Sigma^{-1}-\Sigma^{-1}\Big(\sum_{i=1}^{n}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})^{\top}\Big)\Sigma^{-1}.$ (9)
@@
$\widehat{\Sigma}=\Big(\sum_{i=1}^{n}(T_{i}-1)\Big)^{-1}\Big(\sum_{i=1}^{n}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})^{\top}\Big).$ (10)
As shown, the MLE estimate for $\Sigma$ is obtained. ∎
Appendix D Datasets
HC3: The Human ChatGPT Comparison Corpus (HC3) (Guo et al., 2023) includes comparative responses from human experts and ChatGPT, covering questions from various fields such as open-domain, finance, medicine, law, psychology and Wikipedia. We construct the input by concatenating the Question and Answers together as a single document and label whether it is ChatGPT generated by the source of the answers. We also use the data without Q&A settings and only treat the answer part as a single document.
WikiText: WikiText language modeling dataset (Merity et al., 2016) is a much larger set of verified good and featured articles extracted from Wikipedia compared to WikiSection,we further compare these two dataset (Section D) and show that there is only $\sim 1\%$ potential overlap in topics. We used WikiText-103-v1 collection in specific for experiments. This dataset encompass over 100 million tokens from 29,061 full articles. The dataset is assessible through Huggingface  [^1] https://huggingface.co/datasets/EleutherAI/WikiText_document_level.
Difference between WikiSection and WikiText
[…]
Appendix E Other scores used in this paper
@@
Barzilay and Lapata (2005) is the most recognized entity-based approach. It creates a two-way contingency table for each input document to track the appearance of entities in each sentence. We use Stanford’s CoreNLP to annotate the documents and the implementation provided in the Coheoka library [^2] https://github.com/kigawas/coheoka to obtain the Entity Grid score.
@@
Sheng et al. (2024) introduces BBScore, and also check the main text for a comprehensive comparison between BBScore and BBScoreV2.
Appendix F Human-AI comparison test
Table 5 presents the performance of BBScoreV2 computed with different $\widehat{\Sigma}\in\mathbb{R}^{d}$, while Table 6 shows the performance of the BBScore with various $\widehat{\sigma}\in\mathbb{R}$, where the subscript indicates the dataset used for approximation.
@@ Appendix F Human-AI comparison test
Table 7: Human-AI Task Results with DetectGPT (Mitchell et al., 2023). As a comparison, BBScoreV2 only requires one LLM inference.